# Methods Paper Version 19

### v1

@@ III.1 Naive application
The choice of eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ as SRFs is also convenient from the point of view of many-body applications in combination with the NCSM or other ab initio approaches. These approaches conventionally utilize the M-scheme and thus operate with a huge number of many-body basis functions (Slater determinants) that do not have a definite angular momentum $J$. Therefore, to simplify applications of the Efros method, it is preferable to organize superpositions of these basis functions with a definite $J$ value. The available NCSM codes like MFDn MFDnMaris2010 ; Euro-Par ; Aktulga ; NCSM-NTSE2014 ; Shao and BIGSTICK BIGSTICK provide the NCSM eigenenergies and eigenfunctions with definite $J$ values obtained dynamically during the Hamiltonian diagonalization. So, the most natural and simple way to arrange definite $J$ superpositions of M-scheme Slater determinants is to use the NCSM eigenfunctions corresponding to eigenenergies in the energy range of interest.
Setting $\bar{\beta}_{q}=\beta_{q}^{\textrm{Efn}^{\vphantom{P}}}(\vec{r})$ in Eq. (22) and using the standing-wave functions ${u}(E;\vec{r})$ given by Eq. (19), with the help of Eq. (55) we obtain:
$(E_{q}-E)\langle\beta^{\textrm{Efn}}_{q}|u(E;\vec{r})\rangle=0,\quad q=0,...\,,N-1.$ (58)
Therefore, at any energy $E\neq E_{q}$,
$\langle\beta_{q}^{\textrm{Efn}}|{u}(E;\vec{r})\rangle=0,\quad q=0,...\,,N-1,$ (59)
and the last of these equations with $q=N-1$ results in
$\tan\delta_{l}(E)=\frac{B_{N-1}}{A_{N-1,N-1}}.$ (60)
The resulting phase shifts $\delta_{0}(E)$ in the ${}^{1}S_{0}$ partial wave of $np$ scattering obtained with the Daejeon16 $NN$ interaction Daejeon16 using the oscillator basis with ${\hbar\Omega=30}$ MeV truncated at $N_{{\max}}=16$ are presented in Fig. 1. It is seen that the obtained phase shifts at any $N$ have nothing to do with the exact ones. This issue stems from the structure of the Hamiltonian $H^{tr}$ in Eq. (50) and the use of its eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ as SRFs $\bar{\beta}_{q}(\vec{r})$ which result in producing the wave function ${u}(E;\vec{r})$ as an expansion in a truncated set of oscillator functions $\phi_{n}(\vec{r})$ with ${n\leq\mathcal{\color[rgb]{0,0,0}N}-1}$ with no coupling to the continuum. The wave function components $\phi_{n}(\vec{r})$ with $n=\mathcal{\color[rgb]{0,0,0}N},\mathcal{\color[rgb]{0,0,0}N}+1,...\,,\,\infty$ are also needed since only the infinite sum of oscillator functions in Eq. (51) can describe the continuum spectrum wave function. Therefore, to calculate the contribution of the wave function components $\phi_{n}(\vec{r})$ with $n\geq\mathcal{\color[rgb]{0,0,0}N}$, we obtain an oversimplified Eq. (60) for $\tan\delta_{l}(E)$ which enters Eq. (53) for $a^{as}_{n}(E)$. Equation (60) contains information only about a single eigenstate $\beta_{N-1}^{\textrm{Efn}}(\vec{r})$ of the Hamiltonian $H^{tr}$ and thus cannot describe its complete structure and the interaction $V$ in the system. So, we cannot expect to obtain adequate phase shifts using Eq. (60) in a range of energies $E$.
Figure 1: ${}^{1}S_{0}$ phase shifts $\delta_{0}(E)$ as functions of energy $E$ obtained for various $N$ with $N_{\textrm{max}}=16$ and $\hbar\Omega=30$ MeV using eigenfunctions $\beta_{q}^{\textrm{Efn}}$ as SRFs. Vertical lines indicate the eigenvalues of $H^{tr}$.
One can, however, expect that in this case we obtain at least good phase shifts $\delta_{l}(E_{q})$ at energies $E=E_{q}$ [see Eq. (58)] and can use them for evaluating phase shifts in some energy interval as is done within the SS-HORSE approach AMShirokovHORSE2016 ; BlokhintsevSSHorseC ; AMShirokovHORSE2018 . We see in Fig. 1 that the curves describing phase shifts obtained with different $N$ cross the curve representing the exact phase shifts at energies close to the lowest $E_{q}$ values, however, contrary to the SS-HORSE, calculations of $\tan\delta_{l}(E_{q})$ using Eq. (60) are not accurate and this inaccuracy increases for higher eigenenergies.
$\bar{\beta}^{\rm HO}_{q}(\vec{r})=$ $\phi_{q}(\vec{r}),\qquad q\leq v=N-1,$ (61a)
${\beta}^{\rm HO}_{q}(\vec{r})=$ $\phi_{q}(\vec{r}),\qquad q\leq v-1=N-2.$ (61b)
With the harmonic oscillator SRFs (61), we expect the expansion coefficients of the wave function (51) to be
$d_{n}(E)=\begin{cases}b_{n}(E)+a^{as}_{n}(E),&n=0,...\,,N-2,\\ a^{as}_{n}(E),&n=N-1,...\,,\infty,\end{cases}$ (62)
with $a^{as}_{n}(E)$ given by Eq. (53). The results of the ${}^{1}S_{0}$ phase shift calculations using the oscillator basis with ${\hbar\Omega=30}$ MeV truncated at $N_{{\max}}=16$ are presented in Fig. 2 in comparison with the exact phase shifts for this $NN$ interaction.
Figure 2: ${}^{1}S_{0}$ phase shifts obtained for various $N$ with $N_{\textrm{max}}=16$ and $\hbar\Omega=30$ MeV using the harmonic oscillator SRF. The inset in this and in figures below presenting ${}^{1}S_{0}$ phase shifts shows their low-energy behavior in a larger scale.
In the case of $N=\mathcal{\color[rgb]{0,0,0}N}=9$, the set of SRFs (61) results in the unphysical phase shifts of the same type as the set of eigenfunction SRFs (54) presented in Fig. 1. This is not surprising since both the functions (61a) and (54a) form a complete basis in the Hilbert space spanned by the oscillator functions $\phi_{n}(\vec{r})$ with $n\leq\mathcal{\color[rgb]{0,0,0}N}-1$. Hence, any of the functions $\bar{\beta}^{\rm HO}_{q}(\vec{r})$ can be expressed as a superposition of the functions $\bar{\beta}^{\rm Efn}_{q}(\vec{r})$ and the set of equations (59) guarantees at any energy $E\neq E_{q}$ that
$\langle{\bar{\beta}_{q}^{\textrm{HO}}}|{u}(E;\vec{r})\rangle=0,\quad q=0,...\,,N-1$ (63)
with no coupling to the continuum and results in obtaining expansion coefficients $d_{n}(E)=0$ for all $n=0,$ , $\mathcal{\color[rgb]{0,0,0}N}-1$.
The situation improves dramatically when $N<\mathcal{\color[rgb]{0,0,0}N}$. The phase shifts $\delta_{0}(E)$ demonstrate clear convergence patterns with increasing $N$ and a good description of exact phase shifts is achieved even at relatively small values of $N=5$.
III.2 Proper application by coupling to the continuum
The success of the phase shift calculations with harmonic oscillator SRFs with $N<\mathcal{\color[rgb]{0,0,0}N}$ seems surprising. The problem is that we were using an inconsistent formulation of the problem which disappears in the case of $N<\mathcal{\color[rgb]{0,0,0}N}$ for the oscillator SRFs. In particular, our Hamiltonian (50) was confined to the Hilbert space spanned by the oscillator functions $\phi_{n}(\vec{r})$ with $n\leq\mathcal{\color[rgb]{0,0,0}N}-1$. At the same time, we supposed that our wave function $u(E;\vec{r})$ has an infinite expansion in oscillator functions (51) where, according to Eq. (52), the expansion coefficients $d_{n}(E)=a^{as}_{n}(E)$ at $n\geq\mathcal{\color[rgb]{0,0,0}N}$. The asymptotic expansion coefficients $a^{as}_{n}(E)$ [see Eq. (53)] describing the free motion at energy $E$ and within the HORSE formalism fit the free Schrödinger equation BangHORSE ; Zaitsev1998
$\sum_{n^{\prime}=0}^{\infty}(T^{l}_{nn^{\prime}}-E\delta_{nn^{\prime}\!})\,a^{as}_{n^{\prime}}(E)=0,\qquad n=1,...\,,\,\infty,$ (64)
Therefore, according to our proposition regarding the structure of the wave function expansion coefficients, Eqs. (52)–(53), the Hamiltonian (50) should be extended to infinity in the oscillator space by the kinetic energy matrix elements,
$H=\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle V_{nn^{\prime}}\langle\phi_{n^{\prime}}|+\sum_{n,n^{\prime}=0}^{\infty}|\phi_{n}\rangle T^{l}_{nn^{\prime}}\langle\phi_{n^{\prime}}|\\ \equiv\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle H^{tr}_{nn^{\prime}}\langle\phi_{n^{\prime}}|+\sum_{n,n^{\prime}=\mathcal{\color[rgb]{0,0,0}N}}^{\infty}|\phi_{n}\rangle T^{l}_{nn^{\prime}}\langle\phi_{n^{\prime}}|\\ +|\phi_{\mathcal{\color[rgb]{0,0,0}N}-1}\rangle T^{l}_{\mathcal{\color[rgb]{0,0,0}N}-1,\mathcal{\color[rgb]{0,0,0}N}}\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}}|+|\phi_{\mathcal{\color[rgb]{0,0,0}N}}\rangle T^{l}_{\mathcal{\color[rgb]{0,0,0}N},\mathcal{\color[rgb]{0,0,0}N}-1}\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}-1}|,$ (66)
where the truncated Hamiltonian $H^{tr}$ given by Eq. (50) has the structure conventionally used in variational calculations like NCSM and its matrix elements $H^{tr}_{nn^{\prime}}$ include both potential energy matrix elements $V_{nn^{\prime}}$ and kinetic energy matrix elements $T^{l}_{nn^{\prime}}$ with $n,n^{\prime}=0$, … , $\mathcal{\color[rgb]{0,0,0}N}-1$, the term $\sum_{n,n^{\prime}=\mathcal{\color[rgb]{0,0,0}N}}^{\infty}|\phi_{n}\rangle T^{l}_{nn^{\prime}}\langle\phi_{n^{\prime}}|$ describes the free motion (continuum) of colliding particles, and the last two terms in Eq. (66) describe the coupling between the continuum and the interaction region in the oscillator space.
The role of these interaction region–continuum coupling terms is very important and cannot be neglected. However, in the case of eigenfunction SRFs $\beta_{q}^{\textrm{Efn}}(\vec{r})$ and $\bar{\beta}_{q}^{\textrm{Efn}}(\vec{r})$, these coupling terms were completely suppressed since all matrix elements $\langle\bar{\beta}_{q}^{\textrm{Efn}}|H^{tr}|\phi_{\mathcal{\color[rgb]{0,0,0}N}}\rangle=\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}}|H^{tr}|\beta_{q}^{\textrm{Efn}}\rangle=0$. As a result, we arrived at an unphysical Eq. (60) for the phase shifts. The case of a “complete” set of harmonic oscillator SRFs $\bar{\beta}^{\rm HO}_{q}(\vec{r})$, Eq. (61), with $N=\mathcal{\color[rgb]{0,0,0}N}$ is equivalent to that of the eigenfunction SRFs $\beta_{q}^{\textrm{Efn}}(\vec{r})$ as discussed in the previous subsection. In the expansion of the wave function (51), all coefficients $d_{n}(E)=0$ with $n\leq\mathcal{\color[rgb]{0,0,0}N}-1$ and we have an analog of the scattering by an impenetrable sphere in the coordinate space with the unphysical phase shifts of the same type.
However in the case of an “incomplete” set of harmonic oscillator SRFs $\bar{\beta}^{\rm HO}_{q}(\vec{r})$ with $N<\mathcal{\color[rgb]{0,0,0}N}$, by solving Eq. (33), we obtain a set of functions $g_{n}(E)$ with $n=0$, … , $N-2$ in Eq. (52) and get non-zero expansion coefficients ${d_{n}(E)=a^{as}_{n}(E)}$ with $n\geq N-1$ which mimic the role of the last three terms in Eq. (66) though these terms were not explicitly included in the Hamiltonian of Eq. (50).
So, for the correct application of the Efros formalism with oscillator basis, one should use in the single-channel case the Hamiltonian (66). Regarding the SRFs $\beta_{q}(\vec{r})$, one can either use a set of $N-1$ oscillator functions $\phi_{n}(\vec{r})$ with $n\leq\mathcal{\color[rgb]{0,0,0}N}-1$, eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ of the Hamiltonian $H^{tr}$, or a set of ${N-1}$ linearly-independent superpositions of these functions which may be non-orthogonal. As for the functions $\bar{\beta}_{q}(\vec{r})$, we suggest to use ${N-1}$ functions $\beta_{q}(\vec{r})$ plus an additional function $\bar{\beta}_{N-1}(\vec{r})$ that guarantees the coupling to the continuum. It is most convenient to set it equal to $\phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r})$ which is coupled to a single function $\phi_{\mathcal{\color[rgb]{0,0,0}N}-1}(\vec{r})$ in the interaction region by the analytic matrix element $T^{l}_{\mathcal{\color[rgb]{0,0,0}N}-1,\mathcal{\color[rgb]{0,0,0}N}}$. That is, we choose
$\bar{\beta}_{q}(\vec{r})=\ \begin{cases}\beta_{q}(\vec{r}),&0\leq q\leq N-2,\\ \phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r}),&q=N-1.\end{cases}$ (67)
@@ III.2 Proper application by coupling to the continuum
We note that in the case of $N=\mathcal{\color[rgb]{0,0,0}N}$ (and, generally speaking, with $N>\mathcal{\color[rgb]{0,0,0}N}$) the version of the Efros formalism under discussion is equivalent to the HORSE formalism BangHORSE ; Zaitsev1998 ; Yamani , though these formalisms utilize different equations for calculating phase shifts, wave functions, etc. The advantage of the Efros approach is that it can use a smaller number $N$ of SRFs $\bar{\beta}_{q}(\vec{r})$ than the dimensionality of the truncated Hamiltonian matrix $\mathcal{\color[rgb]{0,0,0}N}$. The HORSE formalism provides exact solutions for the Hamiltonian Eq. (66) and its multichannel extensions including their versions with smoothed interactions [see Eqs. (68) and (75) below]. The Hamiltonian (66) is an approximation since the potential energy is truncated in the oscillator space, however the accuracy of this approximation is improved by increasing the truncation boundary ${\cal N}$. The Efros formalism with $N<{\cal N}~{}$ suggests an approximation of the HORSE and generates solutions which converge to the HORSE solutions with increasing $N$. We shall see below that one can obtain well-converged results with $N<\mathcal{\color[rgb]{0,0,0}N}$ and even with $N\ll\mathcal{\color[rgb]{0,0,0}N}$ in some cases.
Figure 3: ${}^{1}S_{0}$ phase shifts obtained by the HORSE method for various $N_{\textrm{max}}$ with $\hbar\Omega=30$ MeV.
Figure 4: ${}^{1}S_{0}$ phase shifts obtained by the HORSE method for various $\hbar\Omega$ at $N_{\textrm{max}}=24$.
III.3 Convergence of phase shifts and amplitudes
We start with examining the ${}^{1}S_{0}$ phase shifts within the Efros method with the SRF basis set with ${N=\mathcal{\color[rgb]{0,0,0}N}+1}$ which is equivalent to the HORSE method utilizing the same Hamiltonian (66), and investigate the phase shift convergence with $N_{\textrm{max}}$ at various $\hbar\Omega$. Figure 3 shows the convergence with $N_{\textrm{max}}$ at $\hbar\Omega=30$ MeV. It is seen that at $N_{\textrm{max}}=18$, we obtain a very good description of the phase shifts while at $N_{\textrm{max}}=24$, a complete convergence is achieved in a wide range of energies at this particular $\hbar\Omega$ value. Figure 4 shows the change of the phase shifts with $\hbar\Omega$ at $N_{\textrm{max}}=24$. We see that small $\hbar\Omega$ values are less optimal for convergence and henceforth, we use $\hbar\Omega=30$ MeV while studying the $np$ scattering in what follows.
Figure 5: ${}^{1}S_{0}$ phase shifts obtained by the Efros method using a) $N=4$ and $N_{\textrm{max}}=10$, b) $N=4$ and ${N_{\textrm{max}}=14}$, and c) $N=5$ and ${N_{\textrm{max}}=14}$ with both the harmonic oscillator and eigenfunction SRFs, ${\hbar\Omega=30}$ MeV.
Figure 6: ${}^{1}S_{0}$ phase shifts obtained by the Efros method with various $N$ at $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV using the eigenfunction SRFs.
Figure 5 compares the ${}^{1}S_{0}$ phase shifts $\delta_{0}(E)$ obtained via the eigenfunction and harmonic oscillator SRFs for three different truncations, namely $N=4$, ${N_{\textrm{max}}=10}$, $N=4$, ${N_{\textrm{max}}=14}$, and $N=5$, $N_{\textrm{max}}=14$. We observe complicated convergence patterns. On the one hand, it is seen that switching from $N_{\max}=10$ to $N_{\max}=14$ at fixed $N=4$ improves the results with the harmonic oscillator SRFs but the results with the eigenfunction SRFs become even worse. On the other hand, switching from $N=4$ to $N=5$ at fixed $N_{\max}=14$ improves essentially the results with the eigenfunction SRFs but the results with the harmonic oscillator SRFs become slightly worse. However, generally, the results improve with increasing $N$ by more than one unit as well as with essential increase of $N_{\max}$.
In general, it is not obvious whether the harmonic oscillator SRFs or the eigenfunction SRFs perform better for the phase shifts. The lowest-energy results for the eigenfunction SRFs look better at the $N=4$ and $N_{\textrm{max}}=10$ truncations. However, the harmonic oscillator SRFs perform better at other energy intervals such as from $E=5$ MeV to $E=20$ MeV with that same truncation. The eigenfunction SRFs provide better convergence around ${E=45{-}50}$ MeV. Both approaches are comparable in the case of the ${N=5}$, $N_{\textrm{max}}=14$ truncation, though the harmonic oscillator SRFs perform better at the lowest energies up to approximately $E\approx 4$ MeV. From $E=4$ MeV to ${E=15}$ MeV, the eigenfunction SRFs result in a better description of the phase shifts, and the harmonic oscillator SRFs provide slightly better-converged phase shifts from $E=15$ MeV to ${E=35}$ MeV. From $E=35$ MeV to ${E=50}$ MeV, the results are virtually indistinguishable. The two sets of SRFs are comparable in terms of convergence, but in what follows we shall concentrate mostly on the case of the eigenfunction SRFs since they are more viable for many-body applications and suggest a more natural connection to the NCSM.
In Fig. 6 we investigate in more detail the convergence of $\delta_{0}(E)$ obtained by the Efros method with $N$ for $N_{\textrm{max}}=14$. We have noted a very reasonable convergence at $N=5$. However, at larger $N$, namely at $N=8$, the results start deviating significantly from the exact result. The irregularity of convergence is a consequence of the short-range behavior of the $NN$ interaction and a sudden drop-off of its matrix elements $V_{nn^{\prime}}$ at $n,n^{\prime}=\cal N$ due to the sharp truncation at relatively small $N_{\textrm{max}}$.
Figure 7: The effect of smoothing with various parameters $a$ on calculations of the ${}^{1}S_{0}$ phase shifts by the HORSE method with $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV.
Figure 8: ${}^{1}S_{0}$ phase shifts obtained by the HORSE method with various $N_{\textrm{max}}$ at $\hbar\Omega=30$ MeV with smoothing with parameter $a=2.5$.
@@ III.3 Convergence of phase shifts and amplitudes
To address the irregular convergence issue, we smooth the potential truncation as suggested in Refs. HungOne ; HungTwo by replacing in Eq. (66) the potential energy operator $\mathop{\sum}\limits_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle V_{nn^{\prime}}\langle\phi_{n^{\prime}}|$ by
$\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle\tilde{V}_{nn^{\prime}}\langle\phi_{n^{\prime}}|=\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle\sigma_{\mathcal{\color[rgb]{0,0,0}N}-1}^{n}V_{nn^{\prime}}\sigma_{\mathcal{\color[rgb]{0,0,0}N}-1}^{n^{\prime}}\langle\phi_{n^{\prime}}|,$ (68)
@@
$\sigma_{\mathcal{\color[rgb]{0,0,0}N}-1}^{n}=\frac{1-e^{-\left(\!a\frac{n-\mathcal{\color[rgb]{0,0,0}N}}{\mathcal{\color[rgb]{0,0,0}N}}\!\right)^{2}}}{1-e^{-a^{2}}}$ (69)
and $a$ is a dimensionless parameter. The recommended values for the parameter $a$ are from the interval ${a=(2.5,10)}$. Figure 7 demonstrates the effect of smoothing on the phase shift calculations within the HORSE method with $N_{\textrm{max}}=14$. The parameter value of $a=2.5$ seems to be sufficient for our purposes and is henceforth used in our studies of $NN$ scattering.
@@
Figure 9: ${}^{1}S_{0}$ phase shifts obtained by the Efros method with various $N$ at $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV using the eigenfunction SRFs and smoothing with parameter $a=2.5$.
The benefits of smoothing become more apparent in Figs. 8 and 9. We see that smoothing results in excellent convergence of the phase shifts within the HORSE method at $N_{\textrm{max}}=14$ as opposed to the $N_{\textrm{max}}=24$ needed for convergence without smoothing (see Fig. 3). In the case of the Efros method (Fig. 9), we obtain reasonable ${}^{1}S_{0}$ phase shifts with $N=5$ at $N_{{\max}}=14$ while $N=6$ provides very accurate results for $\delta_{0}(E)$ to at least $E=50$ MeV. Contrary to the case of non-smoothed interaction, a further increase of $N$ does not destroy the convergence.
@@
It is important to check whether the proposed approach provides the correct scattering wave function, i. e., we reproduce the coefficients $d_{n}(E)$ in the wave function expansion, Eq. (51). We note that we can expect to obtain correct $d_{n}(E)$ values only in the case when we reproduce, at least with a reasonable accuracy, the scattering phase shifts $\delta_{l}(E)$ which enter Eq. (53) for the expansion coefficients $a^{as}_{n}(E)$. We present in Fig. 10 the convergence with $N$ of the coefficients $d_{n}(E)$ at the energy ${E=0.5}$ MeV where the ${}^{1}S_{0}$ phase shift obtained with smoothing at $N_{\max}=14$ and $\hbar\Omega=30$ MeV is close to the exact. The oscillator amplitudes $d_{n}(E)$ are seen to converge reasonably well. We note that the smoothing results in insignificant changes of the amplitudes $d_{n}(E)$ of the order of the line thickness if the phase shifts obtained with smoothing are very close to those obtained without the smoothing, but the amplitudes $d_{n}(E)$, of course, differ in the case when the smoothing results in the changes of the phase shift values.
Figure 10: Coefficients $d_{n}(E)$ of the wave function expansion (51) at energy ${E=0.5}$ MeV in the ${}^{1}S_{0}$ partial wave obtained with various $N$ at ${N_{\textrm{max}}=14}$ and $\hbar\Omega=30$ MeV with smoothing parameter $a=2.5$ using the eigenfunction SRFs.
@@
In some cases, it may be useful to consider a basis that mixes the harmonic oscillator SRFs and the eigenfunction SRFs. For example, in many-body applications, it may happen that only a few lowest eigenstates are in the energy range of interest and those higher in energy do not significantly contribute to the low-energy phase shifts. In this case, it seems reasonable to introduce a few additional harmonic oscillator SRFs describing the relative motion of colliding fragments in the channel of interest. Therefore, we consider also a couple of sets of non-orthonormal linearly-independent SRFs combining both the harmonic oscillator and the eigenfunctions,
$\bar{\beta}_{q}(\vec{r})=\ \begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...\,,q_{0},\\ \phi_{\mathcal{\color[rgb]{0,0,0}N}-q+q_{0}}(\vec{r}),&q=q_{0}+1,...\,,N-2,\\ \phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r}),&q=N-1,\end{cases}$ (70)
and
$\bar{\beta}_{q}(\vec{r})=\ \begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...\,,q_{0},\\ \phi_{q-q_{0}-1}(\vec{r}),&q=q_{0}+1,...\,,N-2,\\ \phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r}),&q=N-1.\end{cases}$ (71)
Here, $q_{0}+1$ is the number of low-lying eigenfunctions used. We note that for the case $N=\mathcal{\color[rgb]{0,0,0}N}+1$, the HORSE method BangHORSE results are reproduced regardless of the basis combination used. For both Eqs. (70) and (71), the set of SRFs $\{\beta_{q^{\prime}}\},$ $q^{\prime}=0,...\,,N-2$ appearing explicitly in Eq. (19) is the same as the set of SRFs $\{\bar{\beta}_{q}\}$ for $q=0,...\,,N-2$ but it excludes $\bar{\beta}_{N-1}=\phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r})$.
In Fig. 12 we present the ${}^{1}S_{0}$ phase shifts obtained with the SRF set of Eq. (70) with $N_{\textrm{max}}=14$, $\hbar\Omega=30$ MeV, and $N=6$. Vertical lines correspond to the energies of eigenstates of the Hamiltonian $H^{tr}$ at which the phase shifts are well-described AMShirokovHORSE2016 . We see that at least $q_{0}=3$ is needed to get reasonable phase shifts in a wide range of energies while with $q_{0}=4$ the results for the phase shifts are excellent.
Figure 11: ${}^{1}S_{0}$ phase shifts obtained by the Efros method with various $q_{0}$ using the hybrid SRF set given by Eq. (70), $N_{\textrm{max}}=14$, $N=6$, $\hbar\Omega=30$ MeV, and the smoothing parameter $a=2.5$. Vertical lines show the energies of eigenstates of the Hamiltonian $H^{tr}$.
Figure 12: The same as Fig. 12 but using the hybrid SRF set given by Eq. (71) and $N_{\max}=12$.
Figure 12 presents the same results but obtained with the SRF set of Eq. (71) and smaller ${N_{\textrm{max}}=12}$. We see that in this case, we get reasonable results even with $q_{0}=0$ and much more accurate with $q_{0}=1$. Thus it is preferable to use in applications the SRF set of Eq. (71).

### v2

@@ III Single-Channel Scattering
The choice of eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ as SRFs is also convenient from the point of view of many-body applications in combination with the NCSM or other ab initio approaches. These approaches conventionally utilize the M-scheme and thus operate with a huge number of many-body basis functions (Slater determinants) that do not have a definite angular momentum $J$. Therefore, to simplify applications of the adapted Efros method, it is preferable to organize superpositions of these basis functions with a definite $J$ value. The available NCSM codes like MFDn MFDnMaris2010 ; Euro-Par ; Aktulga ; NCSM-NTSE2014 ; Shao and BIGSTICK BIGSTICK provide the NCSM eigenenergies and eigenfunctions with definite $J$ values obtained dynamically during the Hamiltonian diagonalization. So, the most natural and simple way to arrange definite $J$ superpositions of M-scheme Slater determinants is to use the NCSM eigenfunctions corresponding to eigenenergies in the energy range of interest.
@@
We note that in the case of $N=\mathcal{N}$ (and, generally speaking, with $N>\mathcal{N}$) the version of the Efros formalism under discussion is equivalent to the HORSE formalism Zaitsev1998 ; BangHORSE ; Yamani , though these formalisms utilize different equations for calculating phase shifts, wave functions, etc. The advantage of the adapted Efros approach is that it can use a smaller number $N$ of SRFs $\bar{\beta}_{q}(\vec{r})$ than the dimensionality of the truncated Hamiltonian matrix $\mathcal{N}$. The HORSE formalism provides exact solutions for the Hamiltonian Eq. (56) and its multichannel extensions including their versions with smoothed interactions [see Eqs. (61) and (66) below]. The Hamiltonian (56) is an approximation since the potential energy is truncated in the oscillator space, however the accuracy of this approximation is improved by increasing the truncation boundary $\cal N$}. The adapted Efros formalism with $N<{\cal N}$~ suggests an approximation of the HORSE and generates solutions which converge to the HORSE solutions with increasing $N$. We shall see below that one can obtain well-converged results with $N<\mathcal{N}$ and even with $N\ll\mathcal{N}$ in some cases. This feature is especially important because the HORSE method is impractical for many-body problems because it requires the calculation of all eigenstates with a definite $J$ value and $0s$ center-of-mass motion. The number of such eigenstates is very large and is unknown. See Ref. PengYinPLB2024 for a discussion of the number of $1^{-}$ $^{4}\textrm{He}$ eigenstates with $0s$ center-of-mass motion up to energies of 100 MeV above the ground state calculated within NCSM.
In SM (Supplemental, , Section II), we examine first the $^{1}S_{0}$ NNNN phase shifts within the HORSE method utilizing the Hamiltonian (56) and investigate their convergence with $N_{\textrm{max}}$ at various $\hbar\Omega$. We find that the complete convergence is attained at $N_{\textrm{max}}=24$ in a wide range of energies and that $\hbar\Omega=30$ MeV is reasonable for our studies of $np$ scattering in what follows, see Ref. (Supplemental, , Figs. S3 and S4).
@@
To improve convergence, we smooth the potential truncation as suggested in Refs. HungOne ; HungTwo by replacing in Eq. (56) the potential energy operator $\mathop{\sum}\limits_{n,n^{\prime}=0}^{\mathcal{N}-1}|\phi_{n}\rangle V_{nn^{\prime}}\langle\phi_{n^{\prime}}|$ by
$\sum_{n,n^{\prime}=0}^{\mathcal{N}-1}|\phi_{n}\rangle\tilde{V}_{nn^{\prime}}\langle\phi_{n^{\prime}}|=\sum_{n,n^{\prime}=0}^{\mathcal{N}-1}|\phi_{n}\rangle\sigma_{\mathcal{N}-1}^{n}V_{nn^{\prime}}\sigma_{\mathcal{N}-1}^{n^{\prime}}\langle\phi_{n^{\prime}}|,$ (61)
@@
$\sigma_{\mathcal{N}-1}^{n}=\frac{1-e^{-\left(\!a\frac{n-\mathcal{N}}{\mathcal{N}}\!\right)^{\hskip-0.7pt2}}}{1-e^{-a^{2}}}$ (62)
and $a$ is a dimensionless parameter. The recommended values for the parameter $a$ are from the interval a=(2.5,10){a=(2.5,10)}. We find that $a=2.5$ is sufficient for our purposes, see Ref. (Supplemental, , Fig. S5). The smoothing results in an excellent convergence of the phase shifts within the HORSE method at $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV as opposed to $N_{\textrm{max}}=24$ needed for convergence without smoothing at the same $\hbar\Omega$, see Ref. (Supplemental, , Fig. S6) for more details.
As it was pointed out above, the adapted Efros approach suggests an approximation of the HORSE method that seems to be reasonable for many-body applications. We therefore examine the $^{1}S_{0}$ phase shifts within the adapted Efros method utilizing the Hamiltonian (56). As is demonstrated in Ref. (Supplemental, , Fig. S7), the harmonic oscillator SRFs (Supplemental, , Eq. (S5)) and the eigenfunction SRFs are comparable in terms of convergence, but in what follows we shall concentrate mostly on the case of the eigenfunction SRFs since they are more viable for many-body applications and suggest a more natural connection to the NCSM. We also note that smoothing improves convergence within the adapted Efros method, as seen in Ref. (Supplemental, , Figs. S8 and S9).
@@
Figure 1: $^{1}S_{0}$ phase shifts obtained by the adapted Efros method with various NN at $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV using the eigenfunction SRFs and smoothing with parameter $a=2.5$.
@@
In Fig. 1, we investigate the phase shift convergence with $N$ for $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV. We obtain reasonable $^{1}S_{0}$ phase shifts with $N=5$ at $N_{{\max}}=14$ while $N=6$ provides very accurate results for $\delta_{0}(E)$ to at least $E=50$ MeV. We note also (see Ref.  (Supplemental, , Fig. S10)) that the proposed approach provides the correct scattering wave function at the energy E=0.5{E=0.5} MeV where the $^{1}S_{0}$ phase shift obtained with smoothing at $N_{\max}=14$ and $\hbar\Omega=30$ MeV is close to the exact. The smoothing results in insignificant changes of the amplitudes $d_{n}(E)$ of the order of the line thickness if the phase shifts obtained with smoothing are very close to those obtained without the smoothing, but the amplitudes $d_{n}(E)$, of course, differ in the case when the smoothing results in the changes of the phase shift values.
@@
In some cases, it may be useful to consider a basis that mixes the harmonic oscillator SRFs and the eigenfunction SRFs. For example, in many-body applications, it may happen that only a few lowest eigenstates are in the energy range of interest and those higher in energy do not significantly contribute to the low-energy phase shifts. In this case, it seems reasonable to introduce a few additional harmonic oscillator SRFs describing the relative motion of colliding fragments in the channel of interest. We present in Ref. (Supplemental, , Figs. S11 and S12) phase shift convergence with a couple of sets of nonorthonormal linearly independent SRFs (Supplemental, , Eqs. (S8) and (S9)).